# Generalizing Denoising to Non-Equilibrium Structures Improves Equivariant Force Fields

### v1

@@ 3.2.4 Discussion
DeNS can better leverage training data to improve the performance in the following two ways. First, DeNS adds noise to non-equilibrium structures to generate structures with new geometries and therefore naturally achieves data augmentation (Godwin et al., 2022). Second, training DeNS enourages learning a different yet highly correlated interaction. Since we encode forces as inputs and predict the original structures in terms of noise corrections, DeNS enables learning the interaction of transforming forces into structures, which is the inverse of force predictions. As demonstrated in the works of Noisy Nodes (Godwin et al., 2022) and UL2 (Tay et al., 2023), training a single model with multiple correlated objectives to learn different interactions can help the performance on the original task.
@@
DeNS shares similar intuitions to self-supervised learning methods like BERT (Devlin et al., 2019) and MAE (He et al., 2022) and other denoising methods (Vincent et al., 2008; 2010; Godwin et al., 2022; Zaidi et al., 2023) – they remove or corrupt a portion of input data and then learn to predict the original data. Learning to reconstruct data can help learning generalizable representations, and therefore these methods can use the task of reconstruction to improve the performance of downstream tasks. However, unlike those self-supervised learning methods (Devlin et al., 2019; He et al., 2022; Zaidi et al., 2023), DeNS requires force labels for encoding, and therefore, we propose to use DeNS as an auxiliary task optimized along with original tasks and do not follow the previous practice of first pre-training and then fine-tuning. Additionally, we note that before obtaining a single equilibrium structure, we need to run relaxations and generate many intermediate non-equilibrium ones, which is the labeling process as well. We hope that the ability to leverage more from non-equilibrium structures as proposed in this work can encourage researchers to release data containing intermediate non-equilibrium structures in addition to final equilibrium ones. Moreover, we note that DeNS can also be used in fine-tuning. For example, we can first pre-train models on PCQM4Mv2 dataset and then fine-tune them on the smaller MD17 dataset with both the original task and DeNS.
@@ 4.1 OC20 Dataset
Epochs
$\sigma_{\text{high}}$
forces
energy
$12$
$0.125$
19.10
284
$12$
$0.250$
19.11
276
$12$
$0.500$
[…]
$12$
$0.750$
19.43
273
$12$
$1.000$
19.52
274
$20$
$0.250$
18.41
271
$20$
$0.500$
[…]
$20$
$0.750$
18.66
263
$20$
$1.000$
18.66
261
$30$
$0.250$
17.85
265
$30$
$0.500$
[…]
$30$
$0.750$
18.06
255
$30$
$1.000$
18.06
256
Epochs
$p_{\text{DeNS\leavevmode\nobreak\ }}$
forces
energy
$12$
$0.125$
19.64
276
$12$
$0.250$
19.31
273
$12$
$0.500$
[…]
$12$
$0.750$
19.90
281
$20$
$0.250$
18.66
263
$20$
$0.500$
[…]
$20$
$0.750$
18.92
269
$30$
$0.250$
18.06
255
$30$
$0.500$
[…]
$30$
$0.750$
18.24
262
Epochs
$\lambda_{\text{DeNS\leavevmode\nobreak\ }}$
forces
energy
$12$
$5$
19.64
277
$12$
$10$
[…]
$12$
$15$
19.25
275
$20$
$5$
19.02
262
$20$
$10$
18.66
263
$20$
$15$
[…]
$30$
$5$
18.40
260
$30$
$10$
18.06
255
$30$
$15$
[…]
@@
Epochs
forces
energy
# params training time
forces
energy
# params training time
$12$
20.46
285
$83$M 1398
19.32
271
$89$M 1501
$20$
19.78
280
$83$M 2330
18.49
262
$89$M 2501
$30$
19.42
278
$83$M 3495
17.83
255
$89$M 3752
@@
$20$
20.61
290
$52$M 1802
19.07
279
$52$M 1829
Index
forces
energy
1 DeNS
19.32
271
2 Without force encoding
20.16
271
3 $\lambda_{E}=0$ in Eq. 6
19.57
281
4 Fixed $\sigma=0.1$
19.58
283
5 With $\sigma$ encoding
19.66
279
Table 1: Ablation results of EquifomerV2 trained with DeNS on the $2$M subset of the OC20 S2EF dataset. We report mean absolute errors for forces in meV/Å and energy in meV, and lower is better. Errors are averaged over the four validation sub-splits of OC20. The training time is in GPU-hours and measured on V100 GPUs. (a)-(c) We use $\sigma_{\text{high}}=0.5$ for $12$ epochs, $\sigma_{\text{high}}=0.75$ for $20$ and $30$ epochs, $p_{\text{DeNS}}=0.25$, and $\lambda_{\text{DeNS}}=10$ as the default DeNS-related hyper-parameters and vary them to study how they affect the performance. The settings with the best energy-force trade-offs are marked in gray. (d) We train EquiformerV2 and eSCN and compare the results of training with and without DeNS. The results of EquiformerV2 trained with DeNS are from (b). (e) We train EquiformerV2 for $12$ epochs with the best setting in (b) to verify the design choices of DeNS.
@@ 4.1.1 Ablation studies
We use EquiformerV2 (Liao et al., 2023) and S2EF-$2$M split of OC20 to investigate how DeNS-related hyper-parameters affect the performance, compare the results of training with and without DeNS and verify some design choices of DeNS.
Hyperparameters.
In Tables LABEL:tab:sigma, LABEL:tab:prob, LABEL:tab:lambda, we vary the upper bound on standard deviations of Gaussian noise $\sigma_{\text{high}}$, the probability of optimizing DeNS $p_{\text{DeNS\leavevmode\nobreak\ }}$, and the loss coefficient for DeNS $\lambda_{\text{DeNS\leavevmode\nobreak\ }}$ to study how the hyper-parameters of DeNS affect performance. We find that the optimal settings are similar when training for different epochs and have the following observations. First, as we increase $\sigma_{\text{high}}$, force predictions become worse monotonically, while energy predictions improve but saturate at $\sigma_{\text{high}}=0.5$. Second, $p_{\text{DeNS}}=0.5$ and $\lambda_{\text{DeNS}}=10\text{ and }15$ work better than any other values across the three different epochs.
@@
Table LABEL:tab:comparison_with_without_dens summarizes the results of training with and without DeNS. For EquiformerV2, incorporating DeNS as an auxiliary task boosts the performance of energy and force predictions and only increases training time by $7.4\%$ and the number of parameters from $83$M to $89$M. Particularly, EquiformerV2 trained with DeNS for $12$ epochs outperforms EquiformerV2 trained without DeNS for $30$ epochs, saving 2.3×2.3$\times$2.3 × training time. Additionally, we also show that DeNS can be applicable to other equivariant networks like eSCN (Passaro & Zitnick, 2023), and training eSCN with DeNS improves both energy and force MAE while slightly increasing training time by $1.5\%$. The different amounts of increase in training time between EquiformerV2 and eSCN are because they use different modules for noise prediction.
@@
We conduct experiments to verify the design choices of DeNS and summarize the results in Table LABEL:tab:design_choices. All the models follow the best setting of training with DeNS for $12$ epochs in Table LABEL:tab:prob. Comparing Index 1 and Index 2, we show that encoding forces $F(S_{\text{non-eq}})$ in Equations 4 and 6 enables denoising non-equilibrium structures to further improve performance. DeNS without force encoding only results in slightly better force MAE than training without DeNS as in Table LABEL:tab:comparison_with_without_dens. We also compare DeNS with and without force encoding on MD17 in Section 4.3 and find that force encoding is critical. Comparing Index 1 and Index 3, we demonstrate that predicting energy of original structures given corrupted ones can be helpful to the original task. Additionally, we compare the performance of using a fixed $\sigma$ and multi-scale noise, and the comparison between Index 1 and Index 4 shows that multi-scale noise improves both energy and force predictions. Since we sample standard deviation $\sigma$ when using multi-scale noise, we investigate whether we need to encode $\sigma$. The comparison between Index 1 and Index 5 shows that DeNS without $\sigma$ encoding works better, and thus we can use the same approach when we use either a fixed $\sigma$ or multi-scale noise.
@@
eSCN L=6 K=20 (Passaro & Zitnick, 2023) $2.9$ 242.955 993 6242.9559936242.955\,993\,6242.955 993 6 17.090 926 4817.0909264817.090\,926\,4817.090 926 48 $236$ $16.2$ $323$
@@
EquiformerV2 + DeNS ($\lambda_{E}=4$, $160$M) $1.8$ 221 14.2 216 13.4 308
Table 2: OC20 results on S2EF validation and test splits and IS2RE test split when trained on OC20 S2EF-All+MD split. Throughput is reported as the number of structures processed per GPU-second during training and measured on V100 GPUs.
4.1.2 Main results
We train EquiformerV2 ($160$M) with DeNS on the S2EF-All+MD split of OC20. The model follows the same configuration as EquiformerV2 ($153$M) trained without DeNS, and the additional parameters are due to force encoding and one additional block of equivariant graph attention for noise prediction. We report results in Table 2. All test results are computed via the EvalAI evaluation server111eval.ai/web/challenges/challenge-page/712. EquiformerV2 trained with DeNS achieves better S2EF results and comparable IS2RE results, setting the new state-of-the-art results. The improvement is not as significant as that on OC20 S2EF-2M, OC22 (Section 4.2) and MD17 (Section 4.3) datasets since the OC20 S2EF-All+MD training set contains much more structures along relaxation trajectories, making new 3D geometries generated by DeNS less helpful. However, DeNS is valuable because most datasets are not as large as OC20 S2EF-All+MD dataset (about $172$M structures in the training set) but have sizes closer to OC20 S2EF-2M ($2$M structures), OC22 ($8.2$M structures), and MD17 ($950$ structures) datasets.
@@ Results.
We use EquiformerV2 with energy coefficient $\lambda_{E}=4$ and force coefficient $\lambda_{F}=100$ to demonstrate how DeNS can further improve state-of-the-art results and summarize the comparison in Table 3. Compared to EquiformerV2 ($\lambda_{E}=4$, $\lambda_{F}=100$) trained without DeNS, EquiformerV2 trained with DeNS consistently achieves better energy and force MAE across all the S2EF-Total validation and test splits, with $9.6\%$ to $16.7\%$ improvement in energy MAE and $5.1\%$ to $10.4\%$ improvement in force MAE. For IS2RE-Total, EquiformerV2 trained with DeNS achieves the best energy MAE results. The improvement on IS2RE-Total from training with DeNS on only OC22 is greater than that of training on the much larger OC20 and OC22 datasets in previous works. Specifically, training GemNet-OC on OC20 and OC22 datasets (about $134$M + $8.2$M structures) improves IS2RE-Total energy MAE ID by $129$meV and OOD by $50$meV compared to training GemNet-OC on only OC22 dataset ($8.2$M structures). Compared to training without DeNS, training EquiformerV2 with DeNS improves ID by $151$meV and OOD by $167$meV. Thus, training with DeNS clearly improves sample efficiency and performance on OC22.
@@
EquiformerV2 + DeNS ($\lambda_{E}=4,\lambda_{F}=100$) $127$M 391 524 21.13 27.82 233.9 589.9 20.47 30.06 968 1273
@@
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Training time Number of
Model energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces (GPU-hours)$\downarrow$ parameters
SchNet (Schütt et al., 2017) 16.0 58.5 3.5 13.4 3.5 16.9 5.6 28.6 6.9 25.2 8.7 36.9 5.2 24.7 6.1 24.3 - -
DimeNet (Gasteiger et al., 2020) 8.8 21.6 3.4 8.1 2.8 10.0 4.5 16.6 5.3 9.3 5.8 16.2 4.4 9.4 5.0 13.1 - -
PaiNN (Schütt et al., 2021) 6.9 14.7 - - 2.7 9.7 3.9 13.8 5.0 3.3 4.9 8.5 4.1 4.1 4.5 6.0 - -
TorchMD-NET (Thölke & Fabritiis, 2022) 5.3 11.0 2.5 8.5 2.3 4.7 3.3 7.3 3.7 2.6 4.0 5.6 3.2 2.9 4.1 4.1 - -
NequIP ($L_{max}=3$(Batzner et al., 2022) 5.7 8.0 - - 2.2 3.1 3.3 5.6 4.9 1.7 4.6 3.9 4.0 2.0 4.5 3.3 - -
Equiformer ($L_{max}=2$) 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 17 $3.50$M
Equiformer ($L_{max}=3$) 5.3 6.6 2.5 8.1 2.2 2.9 3.2 5.4 4.4 2.0 4.3 3.9 3.7 2.1 4.3 3.4 59 $5.50$M
Equiformer ($L_{max}=2$) + DeNS 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.3 3.9 3.5 1.9 4.2 3.3 19 $4.00$M
Equiformer ($L_{max}=3$) + DeNS 5.0 5.2 2.3 6.1 2.2 2.4 3.2 4.1 3.7 1.6 4.2 3.2 3.4 1.8 4.1 2.9 61 $6.35$M
@@
Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil
Index Attention Layer normalization DeNS energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces
1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3
2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.3 3.9 3.5 1.9 4.2 3.3
3 5.2 7.7 2.4 6.2 2.3 3.9 3.3 6.2 3.8 2.2 4.1 4.7 3.3 2.4 4.2 4.4
4 5.2 6.1 2.4 6.1 2.2 2.9 3.2 5.1 3.7 1.7 4.2 3.9 3.4 2.0 4.2 3.4
5 5.3 9.3 2.4 9.2 2.3 4.5 3.4 8.2 3.7 2.4 4.2 5.5 3.3 2.9 4.2 4.8
6 5.2 7.3 2.4 8.1 2.2 3.4 3.4 6.7 3.7 1.9 4.2 4.4 3.4 2.2 4.2 3.8

### v2

@@ 3.2.4 Discussion
DeNS can better leverage training data to improve the performance in the following two ways. First, DeNS adds noise to non-equilibrium structures to generate structures with new geometries and therefore naturally achieves data augmentation (Godwin et al., 2022). Second, training DeNS enourages learning a different yet highly correlated interaction. Since we encode forces as inputs and predict the original structures in terms of noise corrections, DeNS enables learning the interaction of transforming forces into structures, which is the inverse of predicting forces. As demonstrated in the works of Noisy Nodes (Godwin et al., 2022) and UL2 (Tay et al., 2023), training a single model with multiple correlated objectives to learn different interactions can help the performance on original tasks.
@@
DeNS shares similar intuitions to self-supervised learning methods like BERT (Devlin et al., 2019) and MAE (He et al., 2022) and other denoising methods (Vincent et al., 2008; 2010; Godwin et al., 2022; Zaidi et al., 2023) – they remove or corrupt a portion of input data and then learn to predict the original data. Learning to reconstruct data can help learning generalizable representations, and therefore these methods can use the task of reconstruction to improve the performance on downstream tasks. However, unlike those self-supervised learning methods (Devlin et al., 2019; He et al., 2022; Zaidi et al., 2023), DeNS requires force labels for encoding, and therefore, we propose to use DeNS as an auxiliary task optimized along with original tasks and do not follow the previous practice of first pre-training and then fine-tuning. Additionally, we note that before obtaining a single equilibrium structure, we need to run relaxations and generate many intermediate non-equilibrium ones (Figure 1(a)), which is the labeling process as well. We hope that the ability to leverage more from non-equilibrium structures as proposed in this work can encourage researchers to release data containing intermediate non-equilibrium structures in addition to final equilibrium ones. Moreover, we note that DeNS can also be used in fine-tuning. For example, we can first pre-train models on PCQM4Mv2 dataset and then fine-tune them on the smaller MD17 dataset with both the original task and DeNS.
@@ 4.1 OC20 Dataset
{adjustwidth}
-0.25cm-0.25cm
@@
Number of training time Number of training time
Epochs forces energy parameters (GPU-hours) forces energy parameters (GPU-hours)
$12$ 20.46 285 $83$M 1398 19.09 269 $89$M 1501
$20$ 19.78 280 $83$M 2330 18.58 260 $89$M 2501
$30$ 19.42 278 $83$M 3495 18.02 251 $89$M 3752
@@
$20$ 20.61 290 $52$M 1802 19.14 268 $52$M 1829
Method
Force $\lambda_{E}\neq 0$ Partially corrupted
Index DeNS encoding in Eq. 6 structures forces energy
1 20.46 285
2 19.09 269
3 21.32 278
4 18.87 285
5 19.54 279
Table 1: Ablation results of training with DeNS on OC20 S2EF-2M dataset. We report mean absolute errors for forces in meV/Å and energy in meV, and lower is better. Errors are averaged over the four validation sub-splits of OC20. (a) We train EquiformerV2 and eSCN and compare the results of training with and without DeNS. The training time is measured on V100 GPUs. (b) We train EquiformerV2 for $12$ epochs to verify the design choices of DeNS.
@@ 4.1.1 Ablation Studies
We use OC20 S2EF-2M dataset to compare the results of training with and without DeNS and verify the design choices of DeNS.
@@
Table 1(a) summarizes the results of training with and without DeNS. For EquiformerV2, incorporating DeNS as an auxiliary task boosts the performance on energy and force predictions and only increases training time by $7.4\%$ and the number of parameters from $83$M to $89$M. Particularly, EquiformerV2 trained with DeNS for $12$ epochs outperforms EquiformerV2 trained without DeNS for $30$ epochs and saves 2.3×2.3$\times$2.3 × training time. This suggests that using data augmentation and learning an auxiliary task are more efficient to improve performance compared to simply training for longer. Additionally, we show that DeNS can be applicable to other equivariant networks like eSCN (Passaro & Zitnick, 2023). Training eSCN with DeNS for 20 epochs results in better energy and force MAE compared to EquiformerV2 trained without DeNS for 30 epochs while requiring 1.9×1.9$\times$1.9 × less training time. DeNS slightly increases training time of eSCN by $1.5\%$, and the different amounts of increase in training time between EquiformerV2 and eSCN are because they use different modules for noise prediction.
@@
We train EquiformerV2 for 12 epochs with DeNS to verify the design choices of DeNS and summarize the results in Table 1(b). Comparing Index 2 and Index 3, we show that encoding forces $F(S_{\text{non-eq}})$ in Equations 4 and 6 enables denoising non-equilibrium structures to significantly improve both energy and force MAE. Moreover, DeNS without force encoding (Index 3) results in clearly worse force MAE compared to training without DeNS (Index 1). These results verify our claim that denoising non-equilibrium structures naively can be ill-posed and potentially harmful to performance and that force encoding is critical to leverage the gain of denoising. We note that force encoding only relies on force labels in the training set and does not require any additional data. We also compare DeNS with and without force encoding on MD17 in Section 4.3 and have similar observations. Comparing Index 2 and Index 4, we demonstrate that predicting energy of original structures given corrupted ones can improve energy MAE by $5.6\%$ while slightly increasing force MAE by $1.2\%$. The increase in force MAE is because predicting energy given corrupted structures is equivalent to using $\lambda_{E}\neq 0$ in Equation 6 and implicitly decreases the relative importance of force prediction. Since the decrease in energy MAE is greater than the increase in force MAE, we adopt the practice of predicting energy given corrupted structures. Finally, the comparison between Index 2 and Index 5 shows that partially corrupted structures can further improve the performance gain of DeNS.
4.1.2 Main Results
We train EquiformerV2 ($160$M) with DeNS on OC20 S2EF-All+MD dataset. The model follows the same configuration as EquiformerV2 ($153$M) trained without DeNS, and the additional parameters are due to force encoding and one additional block of equivariant graph attention for noise prediction. We report results in Table 2. All test results are computed via the EvalAI evaluation server111eval.ai/web/challenges/challenge-page/712. EquiformerV2 trained with DeNS achieves new state-of-the-art results on both S2EF and IS2RE tasks. On the S2EF validation split, EquiformerV2 trained with DeNS improves energy MAE by $5$meV and force MAE by $1.0$meV/Å, which is comparable to the gain brought by increasing the size of EquiformerV2 from $31$M to $153$M (energy MAE improved by $5$meV and force MAE improved by $1.3$meV/Å). On the S2EF test split, the improvement in energy and force predictions is smaller, which is probably because of different splits. On the IS2RE test split, training EquiformerV2 with DeNS reduces MAE by $16$meV, achieving similar performance gain of going from eSCN (Passaro & Zitnick, 2023) to EquiformerV2 (Liao et al., 2023) (IS2RE MAE improved by $14$meV) and demonstrating the effectiveness of training with DeNS. We note that the improvement might not be as significant as that on OC20 S2EF-2M (Section 4.1.1), OC22 (Section 4.2) and MD17 (Section 4.3) datasets since the OC20 S2EF-All+MD training set contains much more structures along relaxation trajectories, making new 3D geometries generated by DeNS less helpful. However, DeNS is still valuable because most datasets are not as large as OC20 S2EF-All+MD dataset (about $172$M structures in the training set) but have sizes closer to OC20 S2EF-2M ($2$M structures), OC22 ($8.2$M structures), and MD17 ($950$ structures) datasets.
@@ 4.1.2 Main Results
eSCN L=6 K=20 (Passaro & Zitnick, 2023) $2.9$ 242.955 993 6242.9559936242.955\,993\,6242.955 993 6 17.090 926 4817.0909264817.090\,926\,4817.090 926 48 $228$ $15.6$ $323$
@@
EquiformerV2 + DeNS ($\lambda_{E}=4$, $160$M) $1.8$ 222 14.0 214 13.3 293
Table 2: OC20 results on S2EF validation and test splits and IS2RE test split when trained on OC20 S2EF-All+MD dataset. Throughput is reported as the number of structures processed per GPU-second during training and measured on V100 GPUs.
@@ Results.
We use EquiformerV2 with energy coefficient $\lambda_{E}=4$ and force coefficient $\lambda_{F}=100$ to demonstrate how DeNS can further improve state-of-the-art results and summarize the comparison in Table 3. Compared to EquiformerV2 ($\lambda_{E}=4$, $\lambda_{F}=100$) trained without DeNS, EquiformerV2 trained with DeNS consistently achieves better energy and force MAE across all the S2EF-Total validation and test splits, with $9.6\%$ to $15.3\%$ improvement in energy MAE and $7.1\%$ to $11.7\%$ improvement in force MAE. For IS2RE-Total, EquiformerV2 trained with DeNS achieves the best energy MAE results. The improvement in IS2RE-Total from training with DeNS on only OC22 is greater than that of training on the much larger OC20 and OC22 datasets reported in previous works. Specifically, training GemNet-OC on OC20 and OC22 datasets (about $134$M + $8.2$M structures) improves IS2RE-Total energy MAE ID by $129$meV and OOD by $50$meV compared to training GemNet-OC on only OC22 dataset ($8.2$M structures). Compared to training without DeNS, training EquiformerV2 with DeNS improves ID by $168$meV and OOD by $158$meV. Thus, training with DeNS clearly improves data efficiency and performance on OC22.
@@
EquiformerV2 + DeNS ($\lambda_{E}=4,\lambda_{F}=100$) $127$M 391.6 533.0 20.66 27.11 236.4 590.7 20.04 29.31 951 1282
@@
{adjustwidth}
-2.5cm-2.5cm Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Training time Number of Model energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces (GPU-hours)$\downarrow$ parameters SchNet (Schütt et al., 2017) 16.0 58.5 3.5 13.4 3.5 16.9 5.6 28.6 6.9 25.2 8.7 36.9 5.2 24.7 6.1 24.3 - - DimeNet (Gasteiger et al., 2020) 8.8 21.6 3.4 8.1 2.8 10.0 4.5 16.6 5.3 9.3 5.8 16.2 4.4 9.4 5.0 13.1 - - PaiNN (Schütt et al., 2021) 6.9 14.7 - - 2.7 9.7 3.9 13.8 5.0 3.3 4.9 8.5 4.1 4.1 4.5 6.0 - - TorchMD-NET (Thölke & Fabritiis, 2022) 5.3 11.0 2.5 8.5 2.3 4.7 3.3 7.3 3.7 2.6 4.0 5.6 3.2 2.9 4.1 4.1 - - NequIP ($L_{max}=3$(Batzner et al., 2022) 5.7 8.0 - - 2.2 3.1 3.3 5.6 4.9 1.7 4.6 3.9 4.0 2.0 4.5 3.3 - - Equiformer ($L_{max}=2$) 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 17 $3.50$M Equiformer ($L_{max}=3$) 5.3 6.6 2.5 8.1 2.2 2.9 3.2 5.4 4.4 2.0 4.3 3.9 3.7 2.1 4.3 3.4 59 $5.50$M Equiformer ($L_{max}=2$) + DeNS 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 19 $4.00$M Equiformer ($L_{max}=3$) + DeNS 5.0 5.2 2.3 6.1 2.2 2.4 3.2 4.1 3.7 1.6 4.2 3.2 3.4 1.8 4.1 2.9 61 $6.35$M
@@
{adjustwidth}
-3.5cm-3.5cm Method Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic acid Toluene Uracil Index Attention Layer normalization DeNS energy forces energy forces energy forces energy forces energy forces energy forces energy forces energy forces 1 5.3 7.2 2.2 6.6 2.2 3.1 3.3 5.8 3.7 2.1 4.5 4.1 3.8 2.1 4.3 3.3 2 5.1 5.7 2.3 6.1 2.2 2.6 3.2 4.4 3.7 1.7 4.4 3.7 3.5 1.9 4.2 3.3 3 5.2 7.7 2.4 6.2 2.3 3.9 3.3 6.2 3.8 2.2 4.1 4.7 3.3 2.4 4.2 4.4 4 5.2 6.1 2.4 6.1 2.2 2.9 3.2 5.1 3.7 1.7 4.2 3.9 3.4 2.0 4.2 3.4 5 5.3 9.3 2.4 9.2 2.3 4.5 3.4 8.2 3.7 2.4 4.2 5.5 3.3 2.9 4.2 4.8 6 5.2 7.3 2.4 8.1 2.2 3.4 3.4 6.7 3.7 1.9 4.2 4.4 3.4 2.2 4.2 3.8